OpenStudy (nitroster):

Geometry Check! 3 Q/A. WILL FAN/MEDAL! :)

OpenStudy (anonymous):

AYE SIR!!!

OpenStudy (nitroster):

The last one was too tricky for me. Something that wasn't shown in class :(

OpenStudy (nitroster):

I get stuck at "840 = x(2.1x)" for the last one

OpenStudy (anonymous):

agh the lag!! anyways, 1 and 2 are correct. Now lets go throughout number 3. 3. Looking at the parallel sides, you can assume that, x+40=21+2.1x. Now solve for x. x+40=21+2.1x 40-21=2.1x-x 19=1.1x x=17.27 Btw, I'm completely sure 1 and 2 are correct but number 3 confuzzles me ;-;

OpenStudy (nitroster):

Yeah, me too! Bro, I never looked at it like that! Thx so much! I think thats the closest it can be to corrrect :P AYE MATEYZ!